# On the Training Convergence of Transformers for In-Context Classification

### v1

@@ G.1 Notations
Theorem G.1 (Formal statement of Theorem 4.2)
@@
Proof   The output of the trained transformer is
@@
$\blacksquare$
Remark G.1
We note that Theorem 4.2 requires Assumption 4.2 to hold. For example, we need the covariance $\Lambda$ in training and testing to be the same. A similar consistency requirement of the covariance $\Lambda$ in training and testing had also been observed for in-context linear regression in Zhang et al. [2023a] and for in-context binary classification in the previous section 3.2.
@@
Appendix H Experiment details
@@
For all tasks, we set $d=20$ and we randomly generate a covariance matrix $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$, where $\lambda_{i}=|\hat{\lambda}_{i}|$ and $\hat{\lambda}_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathsf{N}(3,1)$. For each training dataset with different training prompt lengths $N$, and different class numbers $c$, we randomly generate $B$ training samples. Training prompts $P_{\tau},\tau\in[B]$ and their corresponding labels $y_{\tau,\mathsf{query}}$ are generated according to Assumption 4.1. Moreover, we also generate testing datasets. For example, for each testing dataset, we first randomly generate 20 pairs of $(\mu_{j},x_{j,\mathsf{query}},y_{j,\text{prob}}),j\in[20]$, where $(\mu_{j})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{m}_{\Omega}(\Lambda)$, $x_{j,\mathsf{query}}\sim\mathcal{P}^{m}_{x}(\mu_{j},\Lambda)$. $y_{j,\text{prob}}=\mathrm{softmax}(\mu_{j}^{\top}\Lambda^{-1}x_{j,\mathsf{query}})$ are the corresponding probability distributions of the ground truth label $y_{j,\mathsf{query}}$. For each $j$, we generate 100 testing prompts $P_{jk}=(x_{jk,1},y_{jk,1},\dots,x_{jk,M},y_{jk,M},x_{j,\mathsf{query}})$, where $(x_{jk,i},y_{jk,i})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{m}(\mu_{j}\Lambda),j\in[20],k\in[100],i\in[M]$. We denote a model’s output for testing prompts $P_{jk}$ as $\widehat{y}_{jk}$. We calculate its inference error with $\frac{1}{20\times 100}\sum_{j\in[20],k\in[100]}\max_{l\in[c]}\left|\widehat{(}y_{jk})_{l}-(y_{j,\text{prob}})_{l}\right|$, which serves an approximation of the expected total variation distance we defined in (3).
For experiments in Figure 1, we train the single-layer transformers with the sparse-form parameters and structures defined in Section 4. We set the size of the training dataset to $B=10,000$ and set the batch size to 50. We train the transformers using SGD with learning rate $\{0.1,0.5,1\}$ for 10 epochs, and get the best model on each training dataset. Then, we test these trained models on different testing datasets. Each experiment is repeated 10 times with different random seeds. The mean results and standard deviation error bars of these 10 experiments are plotted in Figure 1.
For experiments in Figure 2, the structure of the transformer with full parameters (1-layer, full) is defined as
@@
where $W^{V},W^{KQ}\in\mathbb{R}^{(d+c)\times(d+c)}$ are the parameters for optimization. For the (3-layer, GPT2) model, we use the GPT2 architecture [Radford, 2018] with 64 embedding sizes, 3 layers and 2 heads as implemented by HuggingFace [Wolf et al., 2020], and we use the similar embedding method proposed by Garg et al. [2022]. Note that, for the GPT2 model trained on prompt length $N=100$, it will produce predictions for the labels $\widehat{y}_{i}$, $i\in[101]$. Assume the ground truth labels for the prompt is $y_{i}$, $i\in[101]$, usually, for decoder models like GPT2, we should define the loss function as $\text{Loss}=\sum_{i=1}^{101}\text{CrossEntropyLoss}(\widehat{y}_{i},y_{i})$, which can be viewed to some extent as the model is trained with prompts length ranging from 0 to 100. However, for in-context classification of Gaussian mixtures, longer prompts are significantly more helpful for training models than shorter prompts. For example, considering an extreme case where the classification task involves $c$ classes and the training prompt length is smaller than $c$, then it is nearly impossible for the model to learn to classify $c$ classes Gaussian mixtures in context with these short prompts. Our Theorems 3.1 and 4.1 also clearly demonstrate that the training error is inversely proportional to the prompt length with a scaling of $O(1/N)$. Thus, if we define the loss function for GPT2 as $\text{Loss}=\sum_{i=1}^{101}\text{CrossEntropyLoss}(\widehat{y}_{i},y_{i})$, the underlying short prompts will be detrimental to the model training. Therefore, in our experiments, we define the loss function for our GPT2 model as $\text{Loss}=\sum_{i=41}^{101}\text{CrossEntropyLoss}(\widehat{y}_{i},y_{i})$, which can be viewed to some extent as the model that is trained with prompts length ranging from 40 to 100. For all three transformer models, we set the size of the training dataset to $B=400,000$ and set the batch size to 50. We train the (1-layer, sparse) and (1-layer, full) using Adam with learning rate 0.001 for 5 epochs, and train the GPT2 model using Adam with learning rate 0.0001 for 5 epochs. Each experiment is repeated 3 times with different random seeds. The mean results and standard deviation error bars of these 3 experiments are plotted in Figure 2.

### v3

@@ G.1 Notations
Theorem G.1 (Formal statement of Theorem 4.5).
@@
Proof.
The output of the trained transformer is
@@
Remark G.2.
We note that Theorem 4.5 requires Assumption 4.4 to hold. For example, we need the covariance $\Lambda$ in training and testing to be the same. A similar consistency requirement of the covariance $\Lambda$ in training and testing had also been observed for in-context linear regression in Zhang et al. (2023a) and for in-context binary classification in the previous section 3.2.
@@
Appendix H Additional Experiments
In this section, we provide additional experimental results and the detailed experimental settings.
H.1 Single-layer transformers
(a) $c=10$
(b) $N=80$
(c) log-log axes
(d) log-log axes
We train single-layer transformers for in-context classification of Gaussian mixtures with different numbers of classes $c$, different lengths of training prompts $N$, and test them with different test prompt lengths $M$. The results are reported in Figure 4. We can see from Figure 4 (a,b) that the inference errors decrease as $N$ and $M$ increase, and they increase as $c$ increases. In Figure 4 (c,d), we first fix the training prompt length (test prompt length) to a large number $2000$, and then vary the test prompt length (training prompt length) from 20 to 2000. The results show that, as $M$ and $N$ become sufficiently large, the inference error, which is an approximation of $\mathbb{E}[\Delta(y_{\mathsf{query}},\widehat{y}_{\mathsf{query}})]$ (see Appendix H.2 for detailed definitions), decreases to near-zero. This indicates that the prediction of the trained transformer approaches the Bayes-optimal classifier. All these experimental results corroborate our theoretical claims.
H.2 Experiment Details
@@
For all tasks, we set $d=20$ and we randomly generate a covariance matrix $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$, where $\lambda_{i}=|\hat{\lambda}_{i}|$ and $\hat{\lambda}_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathsf{N}(3,1)$. For each training dataset with different training prompt lengths $N$, and different class numbers $c$, we randomly generate $B$ training samples. Training prompts $P_{\tau},\tau\in[B]$ and their corresponding labels $y_{\tau,\mathsf{query}}$ are generated according to Assumption 4.2. Moreover, we also generate testing datasets. For example, for each testing dataset, we first randomly generate 20 pairs of $(\mu_{j},x_{j,\mathsf{query}},y_{j,\text{prob}}),j\in[20]$, where $(\mu_{j})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{m}_{\Omega}(\Lambda)$, $x_{j,\mathsf{query}}\sim\mathcal{P}^{m}_{x}(\mu_{j},\Lambda)$. $y_{j,\text{prob}}=\mathrm{softmax}(\mu_{j}^{\top}\Lambda^{-1}x_{j,\mathsf{query}})$ are the corresponding probability distributions of the ground truth label $y_{j,\mathsf{query}}$. For each $j$, we generate 100 testing prompts $P_{jk}=(x_{jk,1},y_{jk,1},\dots,x_{jk,M},y_{jk,M},x_{j,\mathsf{query}})$, where $(x_{jk,i},y_{jk,i})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{m}(\mu_{j},\Lambda),j\in[20],k\in[100],i\in[M]$. We denote a model’s output for testing prompts $P_{jk}$ as $\widehat{y}_{jk}$. We calculate its inference error with $\frac{1}{20\times 100}\sum_{j\in[20],k\in[100]}\max_{l\in[c]}\left|\widehat{(}y_{jk})_{l}-(y_{j,\text{prob}})_{l}\right|$, which serves an approximation of the expected total variation distance we defined in (3).
For the ’3-layer’ model, we used the x-transformers library and defined it as an encoder-only transformer with 64 embedding sizes, 3 layers, 2 heads and without positional encoding.
For experiments in Figure 1, we set the size of the training dataset to $B=100,000$ and set the batch size to 50. We train the ’1-layer’ using Adam with learning rate $0.0005$ for 10 epochs, and train the ’3-layer’ using Adam with learning rate $0.0001$ for 5 epochs. Each experiment is repeated 3 times with different random seeds. For experiments in Figure 2, we also set the size of the training dataset to $B=100,000$ and set the batch size to 50. We train the ’1-layer’ using Adam with learning rate $0.001$ for 5 epochs, and train the ’3-layer’ using Adam with learning rate $0.0001$ for 5 epochs. In ’same norm’ and ’same covariance’ settings, pre-training data are sampled according to Assumption 4.2 with a fixed $\Lambda$ that $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$, where $\lambda_{i}=|\hat{\lambda}_{i}|$ and $\hat{\lambda}_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathsf{N}(3,1)$. In ’different norms’ setting, for each $\tau\in[B]$, with probability $\mathbb{P}\left(k=j\right)=1/10,\mu_{\tau,i}\sim\mathsf{N}(k,I_{d}),j=0,1,...,9$, then each Gaussian component is sampled according to $\mathsf{N}(\mu_{\tau,i},\Lambda)$. In (different covariances) setting, we randomly generate $v_{1},v_{2},v_{3}\in\mathbb{R}^{d}$ that half of their elements are 0.1 and the other half elements are 100. Then, we define $\Lambda_{i}=\text{diag}(v_{i}),i=1,2,3$ and generate pre-training data according to Assumption 4.2 with $\Lambda,\Lambda_{1},\Lambda_{2},\Lambda_{3}$. Each experiment is repeated 3 times with different random seeds. For experiments in Figure 3, the structure of the transformer with full parameters ’1-layer, full’ is defined as
@@
where $W^{V},W^{KQ}\in\mathbb{R}^{(d+c)\times(d+c)}$ are the parameters for optimization. For all three transformer models, we set the size of the training dataset to $B=400,000$ and set the batch size to 50. We train the ’1-layer, sparse’ and ’1-layer, full’ using Adam with learning rate 0.001 for 5 epochs, and train the 3-layer transformer model with softmax attention using Adam with learning rate 0.0001 for 5 epochs. Each experiment is repeated 3 times with different random seeds. For experiments in Figure 4, we train the single-layer transformers with the sparse-form parameters and structures defined in Section 4. We set the size of the training dataset to $B=10,000$ and set the batch size to 50. We train the transformers using SGD with learning rate $\{0.1,0.5,1\}$ for 10 epochs, and get the best model on each training dataset. Then, we test these trained models on different testing datasets. Each experiment is repeated 10 times with different random seeds.